OpenStudy (anonymous):

find the total surface area of a hemisphere with a radius of 5 meters.

OpenStudy (anonymous):

the answer at the back of the book shows 235.62m^2 and i know the formula for a sphere is 4 x pie x r^2 so you would just divide by 2 but i keep getting the wrong answer

OpenStudy (jhannybean):

Yep. half a sphere is half of a whole sphere.

OpenStudy (jhannybean):

\[\large \text{whole sphere} = 4\pi r^2 \]\[\large \text{half sphere}= \cfrac{4 \pi r^2}{2} = 2 \pi r^2\]

OpenStudy (jhannybean):

But we're also missing something.

OpenStudy (jhannybean):

\(2 \pi r^2\) only tells us the surface area for the curved part of the sphere. we need the bottom of the circle as well